Please do all four parts for positive feedback. Solve the boundary-value problem RFXZr+Btt11HQX14b4GEZ0DBz2yhAoenJwrPqFQZ for a string of unit length, subject to the given conditions.

  1. u(0,t) = 0, u(1,t) = 0, u(x, 0) = f(x), ut(x, 0) = g(x) f(x) = 1/20 sin πx, g(x) = 0, α = 1/π
  2. u(0,t) = 0, u(1,t) = 0, u(x, 0) = f(x), ut(x, 0) = g(x) f(x) = sin πx cos πx, g(x) = 0, α = 1/π
  3. u(0,t) = 0, u(1,t) = 0, u(x, 0) = f(x), ut(x, 0) = g(x) f(x) = sin πx + 3 sin 2πx − sin 5πx, g(x) = 0, α = 1
  4. g(x) = 0, α = 4, u(0,t) = 0, u(1,t) = 0, u(x, 0) = f(x), ut(x, 0) = g(x)

                                             f(x) = { 2x              0 ≤ x ≤ 1/2

                                                         2(1 − x) 1/2 ≤ x ≤ 1
Answer #1

P76 Itt = a unn - By separation of variables Method U= X(m). THt) rien Xa Ta uan and Чее - Tх ач отрve tя: - т = = -p2, so| And in all four coses qG1o) = g(n) = 0 Their solue to for cq (10) 20 is common for all four- : © Urbino)= & Can.0 + Bninila1: cart) = costit Sintix t3 cos(at) sine in) - Cosoitt) sin(57) @ aza, uhid =foo - fan it oanst 1 [2(6x) of ₂4241 6 » fola 2